OpenStudy (anonymous):

A circular coil of wire 4.22 cm in diameter has a resistance of 4.7 ohms It is located in a magnetic field of 0.149 T directed at right angles to the plane of the coil. The coil is removed from the field in 0.378 s. What is the induced emf in the coil? Assume the magnetic flux is reduced linearly. Answer in units of V

OpenStudy (anonymous):

Not sure what to do with the 4.7 ohms ...but here is the general equation, possible the ohms is to calculate number of turns? otherwise I'm only using one turn N Magnetic flux = change in magnetic field * area M. flux = (0 - 0.149)*pi*.0422^2 = -0.000833 Wb emf = -N * M.flux/ change in time Voltage = -1 * -0.000833 / 0.378 = 0.0022037 Again, that is very small emf, and the induction is made with the number of loops being only 1. If the 4.7 ohms extrapolates to a number of turns, then the voltage will be higher of course. In the very last line for Voltage =, N = 1, but if the number of turns is much greater than that, then the voltage will be multiplied by the number of turns